# AoI-based Scheduling of Correlated Sources for Timely Inference

### v1

@@ Abstract
We investigate a real-time remote inference system where multiple correlated sources transmit observations over a communication channel to a receiver. The receiver utilizes these observations to infer multiple time-varying targets. Due to limited communication resources, the delivered observations may not be fresh. To quantify data freshness, we employ the Age of Information (AoI) metric. To minimize the inference error, we aim to design a signal-agnostic scheduling policy that leverages AoI without requiring knowledge of the actual target values or the source observations. This scheduling problem is a restless multi-armed bandit (RMAB) problem with a non-separable penalty function. Unlike traditional RMABs, the correlation among sources introduces a unique challenge: the penalty function of each source depends on the AoI of other correlated sources, preventing decomposition of the problem into multiple independent Markov Decision Processes (MDPs), a key step in applying traditional RMAB solutions. To address this, we propose a novel approach by approximating the penalty function of each source and establish an analytical bound on the approximation error. We then develop scheduling policies for two scenarios: (i) full knowledge of the penalty functions and (ii) no knowledge of the penalty functions. For the case of known penalty functions, we present an upper bound on the optimality gap of our policy in the asymptotic regime. For the case of unknown penalty functions and signal distributions, we develop an online learning approach that utilizes bandit feedback to learn an online Maximum Gain First (MGF) policy. Simulation results demonstrate the effectiveness of our proposed policies in minimizing inference error and achieving scalability in the number of sources.
@@ I-A Outline and Summary of Contributions
  • To minimize the discounted sum of inference errors for multiple targets, we formulate the problem of scheduling correlated sources over a capacity-limited channel. For the set of all causal and signal-agnostic scheduling policies in which (i) the scheduler makes its decision based on the current and the past information available to the scheduler and (ii) the scheduler does not have access to the realization of the actual processes, we show that the problem can be expressed as minimizing the discounted sum of AoI penalty functions (see Lemma 1 and (13)-(14)).
@@
  • After establishing the approximated function, we develop a Maximum Gain First (MGF) policy (see Algorithm 1) when the scheduler has full knowledge of penalty functions and/or the distribution of targets and source observations. We present an upper bound on the optimality gap of our MGF policy in an asymptotic region (see Theorem 3. The bound suggests that for low correlation settings, our policy yields close to optimal performance for the main problem in an asymptotic region.
@@ I-B Related Works
One of the first works to analyze how AoI values of correlated source observations affect remote inference was [5], in which an information-theoretic analysis was provided to understand the impact of AoI of correlated sources on inference error. However, the authors of [5] did not provide any scheduling policies to optimize inference performance under correlation. In [9], the authors studied the scheduling of correlated sources based on AoI values. This work considers a probabilistic model of correlation between the sources, and then proposed an AoI-based scheduling policy for improving information freshness at the receiver. However, the probabilistic model does not reflect the impact of correlation on inference error or estimation of targets. In [10], the authors developed an AoI-based scheduling policy for minimizing estimation error of correlated Gauss-Markov sources with an LTI system model.
Our current work is related to these prior works on AoI-based scheduling of correlated sources [9, 10]. In particular, this paper generalizes the work of [9, 10] to arbitrary systems, loss functions, and target/source processes. In addition, the prior works [9, 10] consider that the distribution of the signal processes and the penalty functions are known to the scheduler. In this paper, we extend our scheduling policy to the setting, where the distribution of the signal processes and the penalty functions are not available to the scheduler.
@@ II System Model
Consider $M$ sources communicating over a wireless channel to a receiver (Fig. 1). We assume discrete slotted time and at every time slot $t$, each source $m$ observes a time-varying signal $X_{m,t}\in\mathcal{X}_{m}$, where $\mathcal{X}_{m}$ represents the finite set of possible observations from source $m$. A scheduler progressively schedules source observations to the receiver. At each time slot $t$, the receiver uses the received observations to infer $M$ targets $(Y_{1,t},Y_{2,t},\ldots,Y_{M,t})$ of interest. Each target $Y_{m,t}$ is drawn from a finite set $\mathcal{Y}_{m}$ and can be directly inferred from the corresponding source observation $X_{m,t}$. Moreover, the target $Y_{m,t}$ can be correlated with the observation $X_{n,t}$ of any other source $n\neq m$. The results of this paper are derived without assuming a specific correlation model. This ensures the general applicability of our results. An information theoretic interpretation on the impact of the correlation is provided in Sec. IV-A.
@@ II-B Inference Model
Figure 1: A remote inference system with $M$ correlated sources.
@@ III Problem Formulation
where $0<\gamma<1$ is a discount factor, $L(Y_{m,t},a_{m,t})$ is the inference error for the $m$-th target at time $t$, and $\mathrm{V}_{opt}$ is the minimum average inference error. We also consider that the loss function is non-negative and bounded by $B>0$, i.e., we have
$0\leq L(y,a)\leq B$ (4)
for any $y\in\mathcal{Y}_{m}$ and $a\in\mathcal{A}_{m}$.
@@ IV-A An Information-theoretic Interpretation
For $M=2$ sources and $N=1$, we are able to find a low-complexity optimal policy to (13)-(14). According to Theorem 1, the optimal policy is a stationary cyclic policy.
Definition 1 (Stationary Cyclic Policy).
Theorem 1.
$+\sum_{j=0}^{\tau_{2}-1}\sum_{m=1}^{2}\big{(}g_{m}(2+j,1)\big{)}\bigg{)},$ (15)
where $\mathrm{L_{opt}}$ is the optimal objective value of (13)-(14).
Proof.
See Appendix A. ∎
Remark 1.
@@ V Design of Scheduling Policy
In this section, we will explore how to find a low complexity and close to optimal scheduling policy for (13)-(14) (for any values of $M$ and $N$). We also consider that the scheduler knows the penalty function $g_{m}(\cdots)$ for any source $m$. In Section VI, we will design scheduling policy with unknown penalty function.
@@ V-A Function Approximation $f_{m}(\delta_{m})$
Since we have knowledge of the penalty function $g_{m}(\delta_{1},\delta_{2},\ldots,\delta_{M})$, we can evaluate the function $f_{m}(\delta_{m})$, by setting $\delta_{n}=1$ for all $n\neq m$. This effectively evaluates the function under the scenario where all other sources have an AoI of $1$ except source $m$. In Appendix B, we discuss how to compute the function $f_{m}(\delta)$ without using $g_{m}(\cdot,\cdot,\cdot)$.
@@
Our information-theoretic lower bound, $f_{m}(\delta_{m})$, extends the bound in [10] to a significantly broader class of systems, including non-linear and non-Gaussian dynamics. As shown in Appendix B, this bound coincides with the bound in [10, Theorem 2] when specialized to LTI systems with Gaussian processes.
@@ V-B Scheduling Policy
To design a scheduling policy, we first approximate the problem (13)-(14) as follows:
$\mathrm{V}_{f,opt}=$ $\inf_{\pi\in\Pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))\right],$ (20)
@@
where $\mathrm{V}_{f,opt}$ is the optimal objective value of the approximated problem (20)-(21) and $\pi_{f,opt}$ be an optimal policy to (20)-(21) .
@@
$|\mathrm{V}_{{\pi_{f,opt}}}-\mathrm{V}_{opt}|\leq\frac{2}{1-\gamma}\sum_{m=1}^{M}O\bigg{(}\max_{n\neq m}\epsilon_{m,n}^{2}\bigg{)}.$ (22)
@@ V-B1 Lagrangian Relaxation
$\mathrm{V}_{f,opt,\lambda}$
$=\inf_{\pi\in\Pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))+\lambda\pi_{m}(t)\bigg{)}\right]-\frac{\lambda N}{1-\gamma},$ (23)
where $\lambda\in\mathbb{R}$ is the Lagrangian multiplier and $\mathrm{V}_{f,opt,\lambda}$ is the optimal objective value of the relaxed problem (V-B1).
@@
$\max_{\lambda}\mathrm{V}_{f,opt,\lambda},$ (24)
where $\lambda^{*}$ denotes the optimal Langrange multiplier to the problem (24).
@@ V-B2 Gain Index and Threshold Policy
$\inf_{\pi_{m}\in\Pi_{m}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\bigg{(}f_{m}(\Delta_{m}(t))+\lambda\pi_{m}(t)\bigg{)}\right].$ (25)
The value function for each source $m$ is given by
$J_{m,\lambda}(\delta)$
$=f_{m}(\delta)+\min\bigg{\{}\lambda+\gamma J_{m,\lambda}(1),\gamma J_{m,\lambda}(\delta+1)\bigg{\}},$ (26)
where $\delta$ is the current AoI value and $J_{m,\lambda}(\delta)$ is the value function associated with the problem (25). The value function can be obtained by using value iteration algorithm [46].
@@
$J_{m,\lambda}(\delta+1)-J_{m,\lambda}(1)>\frac{\lambda}{\gamma}.$ (27)
@@
Lemma 3 holds due to (V-B2).
Lemma 3 implies that if the difference between $J_{m,\lambda}(\Delta_{m}(t)+1)-J_{m,\lambda}(1)$ exceeds a threshold value $\frac{\lambda}{\gamma}$, then scheduling source $m$ is optimal to the problem (25). However, the policy can not be applied to the main problem due to scheduling constraint. Towards this end, following [7, 25, 37], we define gain index.
@@
For AoI value $\delta$ and Lagrangian multiplier $\lambda$, the gain index is given by
$\alpha_{m,\lambda}(\delta):=J_{m,\lambda}(\delta+1)-J_{m,\lambda}(1).$ (28)
The gain index $\alpha_{m,\lambda}(\delta)$ quantifies the total reduction in the discounted expected sum of inference errors from time $t+1$ to time $\infty$, when action $\pi_{m}(t)=1$ is chosen at time $t$ over $\pi_{m}(t)=0$, where the latter implies $m$-th source is not scheduled at time $t$. This metric enables strategic resource utilization at each time slot to enhance overall system performance. Prior works [7, 25, 37] used action-value function to define the gain index. In [7, 25, 37], source $m$ is scheduled if the difference between two action value $Q_{m,\lambda}(\delta,0)-Q_{m,\lambda}(\delta,1)$ becomes greater than 0, where
$Q_{m,\lambda}(\delta,0)$ $=f_{m}(\delta)+\gamma J_{m,\lambda}(\delta+1)$ (29)
$Q_{m,\lambda}(\delta,1)$ $=f_{m}(\delta)+\lambda+\gamma J_{m,\lambda}(1).$ (30)
Because the instantaneous penalty $f_{m}(\delta)$ incurred at time $t$ is same for whichever decision we take, the policies derived from ($Q_{m,\lambda}(\delta,0)-Q_{m,\lambda}(\delta,1)$) and our gain index-based single source scheduling are equivalent.
For the main problem, our algorithm is provided in Algorithm 1. We call it Maximum Gain First (MGF) Policy. We divide entire duration into episodes, where each episode $k$ contains $T$ time slots. Here, we approximate infinite time with finite time $T$. For discounted penalty problem, the approximation error in considering finite time $T$ is negligible if $T$ is a very large number. At each time $t$, we schedule $N$ sources with highest gain index $\alpha_{m,\lambda_{k}}(\Delta_{m}(t))$. After every episode $k$, we update Lagrange multiplier by using sub-gradient ascent method as follows:
$\lambda_{k+1}$
$=$ $\lambda_{k}+(\theta/k)\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}\mathbf{1}\bigg{(}\alpha_{m,\lambda_{k}}(\Delta_{m}(t))>\frac{\lambda}{\gamma}\bigg{)}-\frac{N}{1-\gamma},$ (31)
where $\mathbf{1}(\cdot)$ is an indicator function, $\theta/k>0$ is the step size, and $T$ is sufficiently large.

### v2

@@ Abstract
We investigate a real-time remote inference system where multiple correlated sources transmit observations over a communication channel to a receiver. The receiver utilizes these observations to infer multiple time-varying targets. Due to limited communication resources, the delivered observations may not be fresh. To quantify data freshness, we employ the Age of Information (AoI) metric. To minimize the inference error, we aim to design a signal-agnostic scheduling policy that leverages AoI without requiring knowledge of the actual target values or the source observations. This scheduling problem is a restless multi-armed bandit (RMAB) problem with a non-separable penalty function. Unlike traditional RMABs, the correlation among sources introduces a unique challenge: the penalty function of each source depends on the AoI of other correlated sources, preventing the problem from decomposing into multiple independent Markov Decision Processes (MDPs), a key step in applying traditional RMAB solutions. To address this, we propose a novel approach that approximates the penalty function for each source and establishes an analytical bound on the approximation error. We then develop scheduling policies for two scenarios: (i) full knowledge of the penalty functions and (ii) no knowledge of the penalty functions. For the case of known penalty functions, we present an upper bound on the optimality gap that highlights the impact of the correlation parameter and the system size. For the case of unknown penalty functions and signal distributions, we develop an online learning approach that utilizes bandit feedback to learn an online Maximum Gain First policy. Simulation results demonstrate the effectiveness of our proposed policies in minimizing inference error and achieving scalability in the number of sources.
@@ I-A Outline and Summary of Contributions
  • To minimize the discounted sum of inference errors for multiple targets, we formulate the problem of scheduling correlated sources over a capacity-limited channel. For the set of all causal and signal-agnostic scheduling policies in which (i) the scheduler makes its decision based on the current and the past information available and (ii) the scheduler does not have access to the realization of the actual processes, we show that the problem can be expressed as minimizing the discounted sum of AoI penalty functions (see Lemma 1 and (12)-(13)).
@@
  • After establishing the approximated function, we develop a Maximum Gain First (MGF) policy (see Algorithm 1) when the scheduler has full knowledge of penalty functions and/or the distribution of targets and source observations. We present an upper bound on the optimality gap of our MGF policy (see Theorem 2). The optimality gap highlights the impact of correlation parameter and the system size.
@@ I-B Related Works
One of the first works to analyze how AoI values of correlated source observations affect remote inference was [5], in which an information-theoretic analysis was provided to understand the impact of AoI of correlated sources on inference error. However, the authors of [5] did not provide any scheduling policies to optimize inference performance under correlation. In [9], the authors studied the scheduling of correlated sources based on AoI values. This work considers a probabilistic model of correlation between the sources, and then proposed an AoI-based scheduling policy for improving information freshness at the receiver. However, the probabilistic model does not reflect the impact of correlation on inference error or estimation of targets. In [10], the authors developed an AoI-based scheduling policy for minimizing estimation error of correlated Gauss-Markov sources with an LTI system model. Our current work is related to these prior works on AoI-based scheduling of correlated sources [9, 10]. In particular, this paper generalizes the work of [9, 10] to arbitrary systems, loss functions, and target/source processes. In addition, the prior works [9, 10] consider that the distribution of the signal processes and the penalty functions are known to the scheduler. In this paper, we extend our scheduling policy to the setting, where the distribution of the signal processes and the penalty functions are not available to the scheduler.
@@ II System Model
Consider $M$ sources communicating over a shared wireless channel to a receiver (Fig. 1). We assume discrete slotted time and at every time slot $t$, each source $m$ observes a time-varying signal $X_{m,t}\in\mathcal{X}_{m}$, where $\mathcal{X}_{m}$ represents the finite set of possible observations from source $m$. A scheduler progressively schedules source observations to the receiver. At each time slot $t$, the receiver uses the received observations to infer $M$ targets $(Y_{1,t},Y_{2,t},\ldots,Y_{M,t})$ of interest. Each target $Y_{m,t}$ is drawn from a finite set $\mathcal{Y}_{m}$ and can be directly inferred from the corresponding source observation $X_{m,t}$. For example, $Y_{m,t}$ can be a function of $X_{m,t}$, such as $Y_{m,t}=X_{m,t}$. Moreover, the target $Y_{m,t}$ can be correlated with the observation $X_{n,t}$ of any other source $n\neq m$. The results of this paper are derived without assuming a specific correlation model. This ensures the general applicability of our results. An information theoretic interpretation on the impact of the correlation is provided in Sec. IV-A.
@@ II-B Inference Model
Figure 1: A remote inference system with $M$ correlated sources, scheduler, shared channel, receiver, and $M$ predictors.
@@ III Problem Formulation
where $0<\gamma<1$ is a discount factor, $L(Y_{m,t},a_{m,t})$ is the inference error for the $m$-th target at time $t$, and $\mathrm{V}_{opt}$ is the minimum average inference error. For the simplicity of theoretical analysis, we consider that the loss function is normalized to $[0,1]$. However, our results can easily be generalized to any bounded loss function.
We also discuss the time-averaged version of the problem in Appendix F.
@@ V Design of Scheduling Policy
In this section, we will explore how to find a low complexity and close to optimal scheduling policy for (12)-(13). We also consider that the scheduler knows the penalty function $g_{m}(\cdots)$ for any source $m$. In Section VI, we will design scheduling policy with unknown penalty function.
@@ V-A Function Approximation $f_{m}(\delta_{m})$
Since we have knowledge of the penalty function $g_{m}(\delta_{1},\delta_{2},\ldots,\delta_{M})$, we can evaluate the function $f_{m}(\delta_{m})$, by setting $\delta_{n}=1$ for all $n\neq m$. This effectively evaluates the function under the scenario where all other sources have an AoI of $1$ except source $m$. In Appendix A, we discuss how to compute the function $f_{m}(\delta)$ without using $g_{m}(\cdot,\cdot,\cdot)$.
@@
Our information-theoretic lower bound, $f_{m}(\delta_{m})$, extends the bound in [10] to a significantly broader class of systems, including non-linear and non-Gaussian dynamics. As shown in Appendix A, this bound coincides with the bound in [10, Theorem 2] when specialized to LTI systems with Gaussian processes.
@@ V-B Scheduling Policy
To design a scheduling policy, we first divide entire duration into episodes, where each episode $k$ contains $T$ time slots. Here, we approximate infinite time with finite time $T$. For discounted penalty problem, the approximation error in considering finite time $T$ is negligible if $T$ is a very large number. Next, we use approximated penalty function. Thus, we reformulated the problem (12)-(13) as follows:
$\mathrm{V}_{f,opt}(T)=$ $\inf_{\pi\in\Pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\gamma^{t}\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))\right],$ (18)
@@
where $T$ is the truncated time horizon, $\mathrm{V}_{f,opt}(T)$ is the optimal objective value of the approximated problem (18)-(19) and $\pi_{f,opt}$ be an optimal policy to (18)-(19).
@@
$|\mathrm{V}_{{\pi_{f,opt}}}(T)-\mathrm{V}_{opt}(T)|\leq\frac{2(1-\gamma^{T})}{1-\gamma}\sum_{m=1}^{M}O\bigg(\max_{n\neq m}\epsilon_{m,n}^{2}\bigg),$ (20)
where $\mathrm{V}_{{\pi_{f,opt}}}(T)$ is the optimal objective value of (18)-(19) and $\mathrm{V}_{opt}(T)$ is the optimal objective value of the problem (12)-(13) with finite horizon $T$.
@@ V-B1 Lagrangian Relaxation
$\mathrm{V}_{f,opt}(\boldsymbol{\lambda},T)$
$=\inf_{\pi\in\Pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\gamma^{t}\bigg(\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))+\lambda_{t}(\pi_{m}(t)-N)\bigg)\right],$ (21)
where $\lambda_{t}\in\mathbb{R}$ is the Lagrangian multiplier, $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\ldots,\lambda_{T})$, and $\mathrm{V}_{f,opt}(\boldsymbol{\lambda},T)$ is the optimal objective value of the relaxed problem (V-B1).
@@
$\max_{\boldsymbol{\lambda}\in\mathbb{R}^{T}}\mathrm{V}_{f,opt,\lambda},$ (22)
where $\boldsymbol{\lambda}^{*}$ denotes the optimal Lagrange value to the problem (22).
@@ V-B2 Gain Index and Threshold Policy
$\inf_{\pi_{m}\in\Pi_{m}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\bigg(f_{m}(\Delta_{m}(t))+\lambda_{t}\pi_{m}(t)\bigg)\right].$ (23)
The optimal value function for each source $m$ at time $t$ is given by
$J_{m,\boldsymbol{\lambda},t}(\delta)$
$=f_{m}(\delta)+\min\bigg\{\lambda_{t}+\gamma J_{m,\boldsymbol{\lambda},t+1}(1),\gamma J_{m,\boldsymbol{\lambda},t+1}(\delta+1)\bigg\},$ (24)
where $\delta$ is the current AoI value and $J_{m,\boldsymbol{\lambda},t}$ is the value function associated with the problem (23). The value function can be obtained by using backward induction method [45].
@@
$J_{m,\boldsymbol{\lambda},t+1}(\delta+1)-J_{m,\boldsymbol{\lambda},t+1}(1)>\frac{\lambda_{t}}{\gamma}.$ (25)
@@
Lemma 3 holds due to (84). ∎
Lemma 3 implies that if the difference between
$J_{m,\boldsymbol{\lambda},t+1}(\delta+1)-J_{m,\boldsymbol{\lambda},t+1}(1)$
exceeds a threshold value $\frac{\lambda_{t}}{\gamma}$, then scheduling source $m$ at time $t$ is optimal to the problem (23). However, the policy can not be applied to the main problem due to scheduling constraint. Towards this end, following [28, 25, 7], we define gain index.
@@
For AoI value $\delta$ and Lagrangian multiplier $\boldsymbol{\lambda}$, the gain index is given by
$\alpha_{m,\boldsymbol{\lambda},t}(\delta):=J_{m,\boldsymbol{\lambda},t+1}(\delta+1)-J_{m,\boldsymbol{\lambda},t+1}(1).$ (26)
The gain index $\alpha_{m,\boldsymbol{\lambda},t}(\delta)$ quantifies the total reduction in the discounted expected sum of inference errors from time $t+1$ to time T1T-1, when action $\pi_{m}(t)=1$ is chosen at time $t$ over $\pi_{m}(t)=0$, where the latter implies $m$-th source is not scheduled at time $t$. This metric enables strategic resource utilization at each time slot to enhance overall system performance. Prior works [7, 25, 28] used action-value function to define the gain index, whereas [47] called it Lagrangian index. In [7, 25, 37, 47], source $m$ is scheduled if the difference between two action value $Q_{m,\boldsymbol{\lambda},t}(\delta,0)-Q_{m,\boldsymbol{\lambda},t}(\delta,1)$ becomes greater than 0, where
$Q_{m,\boldsymbol{\lambda},t}(\delta,0)$ $=f_{m}(\delta)+\gamma J_{m,\boldsymbol{\lambda},t}(\delta+1)$ (27)
$Q_{m,\boldsymbol{\lambda},t}(\delta,1)$ $=f_{m}(\delta)+\lambda_{t}+\gamma J_{m,\boldsymbol{\lambda},t}(1).$ (28)
Because the instantaneous penalty $f_{m}(\delta)$ incurred at time $t$ is same for whichever decision we take, the policies derived from ($Q_{m,\boldsymbol{\lambda},t}(\delta,0)-Q_{m,\boldsymbol{\lambda},t}(\delta,1)$) and our gain index-based single source scheduling are equivalent.
For the main problem, our algorithm is provided in Algorithm 1. We call it Maximum Gain First (MGF) Policy. At each time $t$, we schedule $N$ sources with highest gain index $\alpha_{m,\boldsymbol{\lambda}_{t,k},t}(\Delta_{m}(t))$. If multiple sources have same gain index, we use a tie-breaker index function $\psi_{t,m}(\delta)$ that can rank sources with same gain index value at any time $t$. We can use simple random mixing or Markov random mixing policy to design a tie-breaker index function as discussed in [47].
After every episode $k$, we update Lagrange multiplier by using sub-gradient ascent method as follows:
$\lambda_{t,k+1}=$ $\lambda_{t,k}$
$+$ $\frac{\theta}{k}\bigg(\gamma^{t}\sum_{m=1}^{M}\mathbf{1}\bigg(\alpha_{m,\boldsymbol{\lambda}_{t,k},t}(\Delta_{m}(t))>\frac{\lambda_{t,k}}{\gamma}\bigg)-\frac{N}{1-\gamma}\bigg),$ (29)
where $\mathbf{1}(\cdot)$ is an indicator function and $\theta/k>0$ is step size.